# Diagnosing Hallucination Risk in AI Surgical Decision-Support: A Sequential Framework for Sequential Validation

### v1

@@ 1 Introduction
Large language models (LLMs) are advanced artificial intelligence systems capable of parsing and generating human-like text, demonstrating unprecedented capabilities in healthcare[1, 2, 3]. LLMs offer significant potential to revolutionize healthcare and facilitate clinical workflows[4, 5, 6], by automating data extraction[7], synthesizing medical literature[8], assisting differential diagnoses[9, 10, 11, 12], interpreting radiographic or pathologic reports[1], personalizing patient care by identifying needs and preferences[13, 14], augmenting clinician training[15], and streamlining administrative tasks[16], among others.
@@
Despite their benefits, the use of LLMs still carries inherent risks[17]. Hallucinations, reported in numerous studies[18, 19, 20, 21, 22, 23], are a particular concern in clinical settings. These can manifest as interpretation errors, contextual misalignment, and embedded biases that may jeopardize patient safety and trust[18, 19, 24]. Studies have reported using several frameworks to assess the fidelity mismatch between LLM outputs and healthcare ground truth[18, 25]. This tension between potential benefit and inherent risk highlights the critical need for rigorous, context-aware evaluation frameworks before widespread deployment. However, current hallucination evaluation frameworks account for the severity of errors using taxonomies, the repeatability of LLM outputs, and the potential harms of the errors[18, 25].
@@
To tackle complex reasoning tasks, modern LLMs have employed chain-of-Thoughts (CoT) prompting strategies to enhance LLMs’ ability to parse, analyze, reflect upon, and ultimately solve complex problems[22, 21]. Models incorporating CoT, such as GPT-o1[26] and DeepSeek-r1[27], demonstrate capabilities surpassing conventional instruction-based LLMs, better mimicking human-like logical reasoning for intricate challenges[26, 28, 23]. Much research assessing LLMs in medicine has focused on the potential of serving as diagnostic aids within the physician-AI collaborative paradigm in clinical practice[18]. However, the reliance on the intermediary reasoning steps introduces novel risks, which renders conventional benchmarking metrics inadequate for evaluating the clinical reliability of the outputs from thinking models with CoTs.
@@
On the one hand, current comprehensive benchmarking frameworks for CoT LLMs did not address the clinical healthcare setting[23, 29], such as peri-operative governance. For instance, benchmarking frameworks reported by Jiang et al. only focused on the LLM performances in math, science, optical character recognition (OCR), logic, space-time, and general scenes, neglecting healthcare setting[25].
@@
On the other hand, currently reported evaluation metrics for LLM in healthcare were based on question banks to benchmark LLM performance against medical knowledge mastery within specific disease domains or clinical subspecialties[25, 28, 23, 30, 19]. Such approaches focused on static medical exams or datasets, hence failing to capture the dynamic, interdependent, and often ambiguous nature of real-world clinical reasoning[18, 31]. For instance, Wang et al. only focused on the use of LLM in healthcare conversations, neglecting diagnostics and treatment planning[32]. However, in real-world clinical scenarios, patient symptoms are often non-specific, evidence can conflict, and disease progresses and evolves, while not all auxiliary tests are available.
@@
Therefore, to effectively evaluate the use of LLMs with CoT in real-world peri-operative governance settings, a novel benchmarking evaluation framework is needed to assess the diagnostic precision, quality of clinical action of the LLM-proposed examination and treatment suggestions, the reasoning robustness, output structure, and knowledge alignment[24].
@@
Spinal surgeries present a uniquely compelling focus for such evaluation due to their exceptional clinical complexity and high-stakes decision-making requirements[33, 34, 35, 36, 37, 38]. Spinal disorders represent a massive disease burden worldwide, spanning from common degenerative diseases like degenerative cervical myelopathy, the leading global cause of spinal cord impairment[39, 40], to rather rare and complex diseases like adolescent idiopathic scoliosis, the most common pediatric spinal disorder[41, 42]. Moreover, low back pain alone is a primary driver of disability and the leading reason for medical consultations and abstinence from work[43], with its prevalence and burden increasing with age[44, 45]. This confluence of high clinical prevalence, procedural complexity, and severe consequence profiles establishes spinal surgery as an optimal domain for stress-testing LLMs’ clinical reasoning capabilities and safety alignment in high-risk scenarios.
@@
Finally, we create a transferable deployment foundation that prioritizes interpretability through reasoning chain visualization and safety via real-time hallucination interception. This will establish the auditable standards for LLM integration into high-risk surgical workflows, specifically engineered to mitigate clinical errors at critical multidisciplinary coordination nodes.
@@ 2.3 Disease-Specific Performance Variability
A key finding was DeepSeek-R1’s consistent superiority over DeepSeek-V3 in high-stakes clinical domains, particularly trauma and infection. For trauma, DeepSeek-R1 maintained leading score 45.44 ± 0.96 (Round 1) and 44.11 ± 1.28 (Round 2). Similarly, regarding infection, it achieved 46.44 ± 1.33 and 45.67 ± 1.57 across rounds, demonstrating robust and adaptable inference capabilities despite increasing evaluation complexity.
@@ 3 Discussion
This study demonstrates that while experienced physicians rated the performance of advanced LLMs, our clinician-centered framework reveals a more critical finding: even the most capable models used remain unsuitable for autonomous deployment due to persistent and high-stakes reasoning errors. Superior performance underscores the value of reasoning models, yet no model consistently met the safety thresholds required for unsupervised use in the workflow of spinal surgery. These results align with prior studies highlighting LLMs’ propensity for clinically hazardous inaccuracies[18, 19, 24] and their limited robustness in real-world medical contexts[18, 31]. This reinforces that LLMs are best positioned as decision-support tools under expert clinician oversight, particularly for tasks like differential diagnosis where their outputs align with established paradigms.
@@
The primary merit of this work lies in its novel framework, which leverages deep clinical expertise to move beyond accuracy metrics toward evaluating reasoning fidelity. By stratifying the clinical workflow and embedding high-stakes decision points validated by spinal surgeons, our approach provides a more authentic stress test of model robustness. This approach successfully exposes critical limitations, lacking dynamic guideline alignment, complexity-sensitive reasoning, and clinical safety, that are invisible to simpler benchmarks. Our framework is designed for locating hallucination, which is defined in AI engineering as the production of fluent yet factually inaccurate content with low per-token likelihood, but which is evaluated in the clinical realm solely by its capacity to harm patients through misinformation, irrespective of statistical credibility[18, 24, 46]. Concretely, we capture instances where an autoregressive LLM ventures beyond its training manifold and emits assertions that contradict authoritative medical knowledge or the supplied context[24]. While extant studies have sporadically documented hallucinations within discrete phases of healthcare[24, 17], our methodology provides a comprehensive lens to evaluate this risk across the entire peri-operative continuum by encompassing pre-operative evaluation, intra-operative decision-making, and post-operative surveillance, which remains empirically under-characterized in real clinical populations.
@@
Hallucination management. Mitigating hallucination-induced uncertainty demands a multi-layered, peri-operative governance framework that embeds safeguards at every point where generative models touch spinal care[24]. First, at the data layer, curate continuously updated, subspecialty-specific corpora by integrating operative notes, intra-operative imaging, and post-operative outcomes to narrow the gap between training distribution and real-world variability. Second, at the inference layer, implement retrieval-augmented generation that anchors each token prediction to authoritative ontologies and enforces temporal consistency checks against longitudinal EHR timelines. Third, at the human-AI interface, surgeon-in-the-loop confirmation for any model-generated diagnosis, implant sizing, or complication forecast, supported by visual saliency maps that highlight the imaging voxels or text spans underpinning each recommendation. Fourth, institute post-deployment surveillance: federated error-tracking dashboards that log discrepancies between predicted and observed events across institutions, enabling rapid retraining and bias correction. Finally, cultivate AI-literate surgical teams through scenario-based simulation training and transparent model-cards that clarify limitations and failure modes. Only through this iterative, evidence-driven stewardship can healthcare workers harness generative AI without compromising patient safety or professional accountability.
@@
It is essential to recognize, however, that such architectural safeguards are necessary but insufficient without disciplined model training. Reasoning models are not innate solutions to hallucination; their propensity to fabricate is tightly coupled to their training regimen. Recent evidence demonstrates that models subjected to the full curriculum, supervised fine-tuning followed by reinforcement learning, exhibit markedly fewer hallucinations, whereas those distilled or tuned with reinforcement learning alone succumb to higher fabrication rates[46]. For AI engineers, the implication is unambiguous: adopt the complete training pipeline as a non-negotiable baseline and reserve single-stage shortcuts only for low-risk, well-monitored applications. Continuous post-deployment auditing should then track hallucination incidence across real clinical encounters, enabling iterative retraining that keeps the model’s distribution anchored to authoritative, subspecialty-curated data.
@@
Moreover, while Chain-of-Thought (CoT) prompting enhances the interpretability of AI decision-support systems by elucidating the reasoning process, our findings underscore that it concurrently introduces a distinct class of hallucination risks. The very length and complexity of the reasoning chain, which is intended to mirror clinical cognition, can become a source of error propagation and confabulation. Specifically, (1) Error Accumulation: A minor inaccuracy or unwarranted assumption in an early step of the chain can be amplified through subsequent inferences, leading to a logically structured but fundamentally flawed conclusion[46]. (2) The Illusion of Validity: Elaborate, coherent reasoning can create a persuasive "narrative fallacy," where the sheer volume of plausible-sounding text may obscure underlying factual inaccuracies or gaps in evidence, making it more challenging for clinicians to detect hallucinations than in concise, output-only responses[47]. (3) The Opacity of Complexity: Paradoxically, a lengthy CoT can be as difficult to critically evaluate as a "black box" model, as clinicians may lack the time or resources to verify every logical step in practice[47, 48]. Our sequential validation framework is specifically designed to mitigate these risks by deconstructing the CoT into distinct, verifiable stages, such as initial diagnosis, investigation rationale, and treatment planning. Thereby subjecting each segment of the reasoning process to independent scrutiny. This approach not only helps in localizing where hallucinations originate within the chain but also provides a methodological safeguard against the seductive but potentially misleading coherence of AI-generated rationales.
@@
This study has several important limitations that also delineate the scope of our contribution. As a pilot investigation, the sample size from a single institution is limited, which restricts the statistical power for subgroup analyses and broad generalizability. The exclusive use of a simplified Chinese-language case series, while providing a focused validation context, may introduce confoundings related to variable model pre-training and limits the immediate applicability of findings to other linguistic and clinical settings. Furthermore, the evaluation is restricted to textual reasoning without incorporating radiographic images, which is a critical limitation for a radiographically-driven specialty like spine surgery, impacting the construct validity of the diagnostic assessment. These design choices reflect the pilot nature of this work, which aims to propose and preliminarily validate a framework rather than deliver definitive efficacy evidence.
@@ 4 Conclusion
This framework, detailed for replication in Fig.7, establishes a new reference standard for the translational validation of LLMs. By foregrounding reasoning traceability and workflow-integrated robustness, it positions these tools not as opaque oracles, but as accountable, evidence-bound partners in patient care. We encourage its adoption beyond spinal surgery to ensure the rigorous and safe deployment of AI across all surgical disciplines.
@@ 5.5 Scoring Framework for Clinical Reasoning Assessment
As detailed in Fig 7, clinical reasoning was assessed across six weighted domains: First-visit Diagnosability (max 30 points), Clinical Exam Suggestions (max 30), Treatment Suggestions (max 30), Reasoning Ability (max 60) and Answer Structure (max 10). Domain-specific deductions identified reasoning and generating flaws (e.g., unsorted differentials, omitted key examinations). This granular scoring system provided quantitative proficiency metrics while highlighting targets for educational remediation.

### v2

@@ 1 Introduction
Large language models (LLMs) are advanced artificial intelligence systems capable of parsing and generating human-like text, demonstrating unprecedented capabilities in healthcare[sallam2023, bedi2025, thirunavukarasu2023]. LLMs offer significant potential to revolutionize healthcare and facilitate clinical workflows [iqbal2025, zhang2025b, wang2024], by automating data extraction [lee2025], synthesizing medical literature[croxford2025], assisting differential diagnoses[goh2024a, wu2024, liu2025, jeon2025], interpreting radiographic or pathologic reports [sallam2023], personalizing patient care by identifying needs and preferences [li2025, hunter2024], augmenting clinician training[waldock2024], and streamlining administrative tasks [moor2023], among others.
@@
Despite their benefits, the use of LLMs still carries inherent risks[kwong2024]. Hallucinations, reported in numerous studies[asgari2025, shah2024, goh2024b, wei2022, miao2024, chen2025], are a particular concern in clinical settings. These can manifest as interpretation errors, contextual misalignment, and embedded biases that may jeopardize patient safety and trust[asgari2025, shah2024, yao2025]. Studies have reported using several frameworks to assess the fidelity mismatch between LLM outputs and healthcare ground truth [asgari2025, zhou2024]. This tension between potential benefit and inherent risk highlights the critical need for rigorous, context-aware evaluation frameworks before widespread deployment. However, current hallucination evaluation frameworks mainly account for the severity of errors using taxonomies, the repeatability of LLM outputs, and the potential harms of the errors[asgari2025, zhou2024].
@@
To tackle complex reasoning tasks, modern LLMs have employed chain-of-Thoughts (CoT) prompting strategies to enhance LLMs’ ability to parse, analyze, reflect upon, and ultimately solve complex problems[miao2024, wei2022]. Models incorporating CoT, such as GPT-o1[jaech2024] and DeepSeek-r1[guo2025], demonstrate capabilities surpassing conventional instruction-based LLMs, better mimicking human-like logical reasoning for intricate challenges[jaech2024, mondillo2025, chen2025]. Much research assessing LLMs in medicine has focused on the potential of serving as diagnostic aids within the physician-AI collaborative paradigm in clinical practice [asgari2025]. However, the reliance on the intermediary reasoning steps introduces novel risks, which render conventional benchmarking metrics inadequate for evaluating the clinical reliability of the outputs from thinking models with CoTs.
@@
On the one hand, current comprehensive benchmarking frameworks for CoT LLMs did not address the clinical healthcare setting[chen2025, zhang2025a], such as peri-operative governance. For instance, benchmarking frameworks reported by Jiang et al. only focused on the LLM performances in math, science, optical character recognition (OCR), logic, space-time, and general scenes, neglecting healthcare setting [zhou2024].
@@
On the other hand, currently reported evaluation metrics for LLM in healthcare were based on question banks to benchmark LLM performance against medical knowledge mastery within specific disease domains or clinical subspecialties[zhou2024, mondillo2025, chen2025, li2024, shah2024]. Such approaches focused on static medical exams or datasets, hence failing to capture the dynamic, interdependent, and often ambiguous nature of real-world clinical reasoning[asgari2025, wang2025c]. For instance, Wang et al. only focused on the use of LLM in healthcare conversations, neglecting diagnostics and treatment planning[wang2025b]. However, in real-world clinical scenarios, patient symptoms are often non-specific, evidence can conflict, and disease progresses and evolves, while not all auxiliary tests are available.
@@
Therefore, to effectively evaluate the use of LLMs with CoT in real-world peri-operative governance settings, a novel benchmarking evaluation framework is needed to assess the diagnostic precision, quality of clinical action of the LLM-proposed examination and treatment suggestions, the reasoning robustness, output structure, and knowledge alignment[yao2025].
@@
Spinal surgeries present a uniquely compelling focus for such evaluation due to their exceptional clinical complexity and high-stakes decision-making requirements[zhang2019, yoo2019, choi2017, cheriyan2015, hanada2021, mallepally2020]. Spinal disorders represent a massive disease burden worldwide, spanning from common degenerative diseases like degenerative cervical myelopathy, the leading global cause of spinal cord impairment[nouri2015, witiw2017], to rather rare and complex diseases like adolescent idiopathic scoliosis, the most common pediatric spinal disorder[muhly2016, martin2014]. Moreover, low back pain alone is a primary driver of disability and the leading reason for medical consultations and abstinence from work[hart1995], with its prevalence and burden increasing with age[gbd2018, hoy2012]. This confluence of high clinical prevalence, procedural complexity, and severe consequence profiles establishes spinal surgery as an optimal domain for stress-testing LLMs’ clinical reasoning capabilities and safety alignment in high-risk scenarios.
@@
Finally, we create a transferable deployment foundation that prioritizes interpretability through reasoning chain visualization and safety through real-time hallucination interception. This will establish standards for the integration of LLM into high-risk surgical workflows, specifically engineered to mitigate clinical errors at critical multidisciplinary coordination nodes.
@@ 2.3 Disease-Specific Performance Variability
A key finding was DeepSeek-R1’s consistent superiority over DeepSeek-V3 in high-stakes clinical domains, particularly trauma and infection. For trauma, DeepSeek-R1 maintained a leading score of 45.44 ± 0.96 (Round 1) and 44.11 ± 1.28 (Round 2). Similarly, regarding infection, it achieved 46.44 ± 1.33 and 45.67 ± 1.57 across rounds, demonstrating robust and adaptable inference capabilities despite increasing evaluation complexity.
@@ 3 Discussion
This study demonstrates that while experienced physicians rated the performance of advanced LLMs, our clinician-centered framework reveals a more critical finding: even the most capable models used remain unsuitable for autonomous deployment due to persistent and high-stakes reasoning errors. Superior performance underscores the value of reasoning models, yet no model consistently met the safety thresholds required for unsupervised use in the workflow of spinal surgery. These results align with prior studies highlighting LLMs’ propensity for clinically hazardous inaccuracies[asgari2025, shah2024, yao2025] and their limited robustness in real-world medical contexts[asgari2025, wang2025c]. This reinforces that LLMs are best positioned as decision-support tools under expert clinician oversight, particularly for tasks like differential diagnosis where their outputs align with established paradigms.
@@
The primary merit of this work lies in its novel framework, which leverages deep clinical expertise to move beyond accuracy metrics toward evaluating reasoning fidelity. By stratifying the clinical workflow and embedding high-stakes decision points validated by spinal surgeons, our approach provides a more authentic stress test of model robustness. This approach successfully exposes critical limitations, lacking dynamic guideline alignment, complexity-sensitive reasoning, and clinical safety, that are invisible to simpler benchmarks. Our framework is designed for locating hallucination, which is defined in AI engineering as the production of fluent yet factually inaccurate content with low per-token likelihood, but which is evaluated in the clinical realm solely by its capacity to harm patients through misinformation, irrespective of statistical credibility[asgari2025, yao2025, lanham2023]. Concretely, we capture instances where an autoregressive LLM ventures beyond its training manifold and emits assertions that contradict authoritative medical knowledge or the supplied context[yao2025]. While extant studies have sporadically documented hallucinations within discrete phases of healthcare[yao2025, kwong2024], our methodology provides a comprehensive lens to evaluate this risk across the entire peri-operative continuum by encompassing pre-operative evaluation, intra-operative decision-making, and post-operative surveillance, which remains empirically under-characterized in real clinical populations.
@@
Hallucination management mitigating hallucination-induced uncertainty demands a multi-layered, peri-operative governance framework that embeds safeguards at every point where generative models touch spinal care[yao2025]. First, at the data layer, curate continuously updated, subspecialty-specific corpora by integrating operative notes, intra-operative imaging, and post-operative outcomes to narrow the gap between training distribution and real-world variability. Second, at the inference layer, implement retrieval-augmented generation that anchors each token prediction to authoritative ontologies and enforces temporal consistency checks against longitudinal EHR timelines. Third, at the human-AI interface, surgeon-in-the-loop confirmation for any model-generated diagnosis, implant sizing, or complication forecast, supported by visual saliency maps that highlight the imaging voxels or text spans underpinning each recommendation. Fourth, institute post-deployment surveillance: federated error-tracking dashboards that log discrepancies between predicted and observed events across institutions, enabling rapid retraining and bias correction. Finally, cultivate AI-literate surgical teams through scenario-based simulation training and transparent model-cards that clarify limitations and failure modes. Only through this iterative, evidence-driven stewardship can healthcare workers harness generative AI without compromising patient safety or professional accountability.
@@
It is essential to recognize, however, that such architectural safeguards are necessary but insufficient without disciplined model training. Reasoning models are not innate solutions to hallucination; their propensity to fabricate is tightly coupled to their training regimen. Recent evidence demonstrates that models subjected to the full curriculum, supervised fine-tuning followed by reinforcement learning, exhibit markedly fewer hallucinations, whereas those distilled or tuned with reinforcement learning alone succumb to higher fabrication rates[lanham2023]. For AI engineers, the implication is unambiguous: adopt the complete training pipeline as a non-negotiable baseline and reserve single-stage shortcuts only for low-risk, well-monitored applications. Continuous post-deployment auditing should then track hallucination incidence across real clinical encounters, enabling iterative retraining that keeps the model’s distribution anchored to authoritative, subspecialty-curated data.
@@
Moreover, while Chain-of-Thought (CoT) prompting enhances the interpretability of AI decision-support systems by elucidating the reasoning process, our findings underscore that it concurrently introduces a distinct class of hallucination risks. The very length and complexity of the reasoning chain, which is intended to mirror clinical cognition, can become a source of error propagation and confabulation. Specifically, (1) Error Accumulation: A minor inaccuracy or unwarranted assumption in an early step of the chain can be amplified through subsequent inferences, leading to a logically structured but fundamentally flawed conclusion[lanham2023]. (2) The Illusion of Validity: Elaborate, coherent reasoning can create a persuasive "narrative fallacy," where the sheer volume of plausible-sounding text may obscure underlying factual inaccuracies or gaps in evidence, making it more challenging for clinicians to detect hallucinations than in concise, output-only responses[barez2025]. (3) The Opacity of Complexity: Paradoxically, a lengthy CoT can be as difficult to critically evaluate as a "black box" model, as clinicians may lack the time or resources to verify every logical step in practice[barez2025, shojaee2025]. Our sequential validation framework is specifically designed to mitigate these risks by deconstructing the CoT into distinct, verifiable stages, such as initial diagnosis, investigation rationale, and treatment planning. Thereby, subjecting each segment of the reasoning process to independent scrutiny. This approach not only helps in localizing where hallucinations originate within the chain but also provides a methodological safeguard against the seductive but potentially misleading coherence of AI-generated rationales.
@@
This study has several important limitations that also delineate the scope of our contribution. As a pilot investigation, the sample size from a single institution is limited, which restricts the statistical power for subgroup analyses and broad generalizability. The exclusive use of a simplified Chinese-language case series, while providing a focused validation context, may introduce confoundings related to variable model pre-training and limit the immediate applicability of findings to other linguistic and clinical settings. Furthermore, the evaluation is restricted to textual reasoning without incorporating radiographic images, which is a critical limitation for a radiographically-driven specialty like spine surgery, impacting the construct validity of the diagnostic assessment. These design choices reflect the pilot nature of this work, which aims to propose and preliminarily validate a framework rather than deliver definitive efficacy evidence.
@@ 4 Conclusion
This framework, detailed for replication in Fig 7, establishes a new reference standard for the translational validation of LLMs. By foregrounding reasoning traceability and workflow-integrated robustness, it positions these tools not as opaque oracles, but as accountable, evidence-bound partners in patient care. We encourage its adoption beyond spinal surgery to ensure the rigorous and safe deployment of AI across all surgical disciplines.
@@ 5.5 Scoring Framework for Clinical Reasoning Assessment
As detailed in Fig 7, clinical reasoning was assessed across five weighted domains: First-visit Diagnosability (max 30 points), Clinical Exam Suggestions (max 30), Treatment Suggestions (max 30), Reasoning Ability (max 60) and Answer Structure (max 10). Domain-specific deductions identified reasoning and generating flaws (e.g., unsorted differentials, omitted key examinations). This granular scoring system provided quantitative proficiency metrics while highlighting targets for educational remediation.